OpenStudy (anonymous):

the pressure p volume v and temp t of a gas are related by pv=kt where k is a constant. the total differential dT equals

ganeshie8 (ganeshie8):

\(T = \frac{PV}{k}\) \(dT = \frac{\partial T }{\partial P}dP + \frac{\partial T }{\partial V}dV\)

OpenStudy (anonymous):

thx i got this V dp + p dv. is that right?

ganeshie8 (ganeshie8):

Correct ! except for the missing 1/k